OpenStudy (anonymous):

a coin is loaded so that the probabilityof getting heads 3/4. if the coin is flipped twice,what is the probability of gettiing heads twiice

OpenStudy (anonymous):

2/4 i think

OpenStudy (anonymous):

This is the same kind of problem as your other one. The probability of getting heads the first time is?

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

No, the question specifically says that the coin is loaded so it's 3/4

OpenStudy (anonymous):

o

OpenStudy (anonymous):

And on the second flip you still have another 3/4 chance of getting heads.

OpenStudy (anonymous):

6/8

OpenStudy (anonymous):

So the probability of getting both is the probability of the first, times the probability of the second.

OpenStudy (anonymous):

9/16

OpenStudy (anonymous):

Correct. Anytime you have 2 different events that both must occur you have to multiply the probability for each event together.

OpenStudy (anonymous):

Same thing for the letters/numbers problem you were asking about before. It would be 3/6 chance of a number being even 1/3 chance of a letter being a So probability of both would be 3/6 times 1/3 = 3/18 = 1/6